For point charge -2.6 µC and point charge 7.7 µC located at the same positions as in the previous question (5 m and 4 m, respectively), determine the direction of the net electric field E at the origin.

1µC = 10-6C

Your answer should be an integer, do not include the unit.

Hint: The direction is indicated by an angle measured in degrees counterclockwise from the positive x axis.
Answer #1

Electric field, E = kq/r^2

Ea = (9 x 10^9 x 2.6 x 10^-6)/5^2 = 936 N/C

Eb = (9 x 10^9 x 7.7 x 10^-6)/4^2 = 4331.25 N/C

Direction, theta = arctan (Eb/Ea) = 77.81 degree

Comment in case any doubt please rate my answer....